A neutron in a reactor makes an elastic headon collision with the nucleus of an atom initially at rest. Assume: The mass of the atomic nucleus is about 11.3 the mass of the neutron. What fraction of the neutron’s kinetic energy is transferred to the atomic nucleus?


1
Expert's answer
2021-12-12T16:41:29-0500

m = mass of neutron

M = mass of nucleus = 11.3 m

vi = initial velocity of neutron before collision = ?

Vi = initial velocity of nucleus before collision = 0 m/s

vf = final velocity of neutron after collision

Vf = final velocity of nucleus after collision

using conservation of momentum

"mv_i + MV_i = mv_f + MV_f \\

(m) v_i + (11.3 m) (0) = m v_f + (11.3 m) V_f \\

v_i = v_f + (11.3) V_f \\

v_i - v_f = (11.3) V_f \;\;\; eq-1"

using conservation of kinetic energy

"m v^2_i + M V^2_i = m v^2_f + M V^2_f \\

(m) v^2_i + (11.3 m) (0)^2 = m v^2_f + (11.3 m) V^2_f \\

v^2_i = v^2_f + (11.3) V^2_f \\

v^2_i - v^2_f = (11.3) V^2_f \\

( v_i - v_f )( v_i + v_f ) = (11.3) V^2_f \\

(13.7) V_f ( v_i + v_f ) = (11.3) V^2_f \\

v_i + v_f = V_f \;\;\;\; eq-2"

adding eq-1 and eq-2

"( v_i - v_f ) + ( v_i - v_f ) = (11.3) V_f + V_f \\

2 v_i = 12.3 V_f \\

v_i = 6.15 V_f"

fraction of energy transferred is given as

"fraction = \frac{(0.5) MV^2_f }{(0.5) m v^2_i } \\

fraction = \frac{M V^2_f }{m v^2_i } \\

fraction = \frac{(11.3 m) V^2_f }{m (6.15 V_f )^2 } \\

fraction = \frac{11.3}{(6.15)^2} = 0.298 ≈ 0.3"